OpenStudy (anonymous):

what is the midpoint of a segment with the endpoints m(-18,13), n(44, -13) ? If you answer it, I'll give you a fan and medal :)

OpenStudy (anonymous):

the midpoint of a segment with the endpoints m(x,y),n(a,b) is (i,j) i=(x+a)/2 j=(y+b)/2

OpenStudy (anonymous):

I don't get it...can you give me the answer? lol.. I promise I'll get it once you give me it xD

OpenStudy (anonymous):

(-18+44)/2=13 (13+(-13))/2=0 (13,0)

OpenStudy (anonymous):

oh wow. I actually had that answer. Thanks :)